OpenStudy (anonymous):

If f(1) = 5 and f(x+1) =2f(x) then the value of f(7) is

OpenStudy (anonymous):

\[f(1)=5\] \[f(2)=2\times 5=10\] \[f(3)=2\times 10=20\]

OpenStudy (anonymous):

and in general \[f(x)=5\times 2^{x-1}\]

OpenStudy (anonymous):

making \[f(7)=5\times 2^6\]